# Interpretability and Generalization Bounds for Learning Spatial Physics

### v1

@@ A.2 Concrete Example of Error for 3 point stencil
For a given stencil, the above procedure can be executed algebraically with fewer assumptions. As above, we consider the limit of infinitely many samples of functions $f,u$ with $c\sim C$ and integrating over $x\in[0,1]$.
$\mathcal{L}=\mathbb{E}_{c}\left[\int_{0}^{1}\left(wFD(u,\Delta x)-ku^{\prime\prime}\right)^{2}\mathrm{d}x\right].$ (29)
Thus, given a training set of $u$ and $f$ generated by sampling coefficients $c_{n}$, the inner value of (7) is
$\frac{w}{\Delta x^{2}}\left(\left(\sum_{n=0}^{p+2}c_{n}(x-\Delta x)^{n}\right)-2\left(\sum_{n=0}^{p+2}c_{n}x^{n}\right)+\left(\sum_{n=0}^{p+2}c_{n}(x+\Delta x)^{n}\right)\right)\\ -k\sum_{n=2}^{p+2}n(n-1)c_{n}x^{n-2}$ (30)
meaning the interior of the MSE loss
$\mathcal{L}=\mathbb{E}_{c}\left[\int_{0}^{1}\left(\frac{w}{\Delta x^{2}}\sum_{n=2}^{p+2}c_{n}\left((x-\Delta x)^{n}-2x^{n}+(x+\Delta x)^{n}\right)\\ -k\sum_{n=2}^{p+2}n(n-1)c_{n}x^{n-2}\right)^{2}\mathrm{dx}\right].$ (31)
For $p=0$ and $p=1$, simplifying the finite difference stencil results in an equation where $w=k$ within the squared error term.
But when we consider $p=2$ with quadratic $f$ and quartic $u$, there is not full cancellation:
[…]
@@
which satisfies the general result.
A.3 Proof of Theorem 2
@@
A.4 Corollary to Theorem 2: A Loose Error Bound
A loose error bound naturally follows from this:
Corollary 1.
Assuming that the basis set has $p$ entries, the error $\|{\bm{W}}^{*}-{\bm{A}}\|_{F}/\|{\bm{A}}\|_{F}$ is dependent on the size of the basis set,
$\frac{\|{\bm{W}}^{*}-{\bm{A}}\|_{F}}{\|{\bm{A}}\|_{F}}\leq\sqrt{N_{\text{grid}}-p}\left(\frac{\|{\bm{W}}^{0}\|_{F}}{\|{\bm{A}}\|_{F}}+1\right)$ (42)
Proof.
Note that, ${\bm{W}}^{*}-{\bm{A}}={\bm{W}}^{0}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})-{\bm{A}}({\bm{I}}-{\bm{U}}{\bm{U}}^{T})=({\bm{W}}^{0}-{\bm{A}})({\bm{I}}-{\bm{U}}{\bm{U}}^{T})$. The Frobenious norm satisfies the sub-multiplicative property, meaning $\|({\bm{W}}^{0}-{\bm{A}})({\bm{I}}-{\bm{U}}{\bm{U}}^{T})\|_{F}\leq\|{\bm{W}}^{0}-{\bm{A}}\|_{F}\|{\bm{I}}-{\bm{U}}{\bm{U}}^{T}\|_{F}$. The matrix ${\bm{U}}{\bm{U}}^{T}$ is rank $p$. The complementary matrix has rank $N_{\text{grid}}-p$. The F-norm for a rank $r$ orthogonal matrix is $\sqrt{r}$ by definition. Using the additive distribution of the F-norm, $\|{\bm{W}}^{0}-{\bm{A}}\|=\|{\bm{W}}^{0}\|+\|{\bm{A}}\|$. Assembling these gives the stated result. ∎
A.5 Corollary to Theorem 1: A Tight Error Bound
If the Green’s function is known, the basis ${\bm{B}}$ and analytical solution ${\bm{A}}$ can be computed to give a much tighter estimate, as is shown in Fig.2. The procedure is as follows. Compute a matrix of basis functions evaluated at each point in the domain,
@@ A.5 Corollary to Theorem 1: A Tight Error Bound
${\bm{A}}_{ij}=\int_{0}^{1}\psi_{i}(s)G(s,x_{j})\mathrm{d}x$ (45)
where $\psi_{i}$ is a piecewise linear or a piecewise constant basis function centered on $x_{i}$. Then, the error starting at any initial matrix ${\bm{W}}^{0}$ can be directly computed as
$\|{\bm{W}}^{*}-{\bm{A}}\|_{F}\approx\|{\bm{W}}^{0}-{\bm{A}}{\bm{U}}{\bm{D}}{\bm{V}}^{T}\|_{F}$ (46)
where the computation for ${\bm{A}}$ has error turns from numerical roundoff and the arbitrariness of the choice of discretization basis $\psi(x)$.
@@ Appendix B Procedure Details
To initially rule out errors from software drivers or numerical precision, we replicated the linear models and deep models in many hardware and software variations: Tinygrad, JAX, and PyTorch, on CPU, NVIDIA 4070 GPU, and Apple Silicon GPU, and varied precision in float32 and float64 on CPU. The ultimate experiments ran in float32 on GPUs on a Macbook Pro M3 Max. A single training run for any of the experiments take between 30 seconds to a few minutes on consumer hardware. A full sweep can take 20 minutes. All replications with five seeds takes about 48 hours of compute time on a single commodity laptop. Code scripts to replicate the experiments is included in the supplemental material.
Tuning of hyperparameters of hidden size and learning rate schedule was guided by achieving below $10^{-14}$ train MSE on a $p\geq 5$ function space. The most important aspect in each model was removing weight decay and having a schedule that terminates at a learning rate of $10^{-5}-10^{-6}$.
Finite difference experiments were performed using pysindy (de Silva et al., 2020; Kaptanoglu et al., 2022), running on CPU in float64 precision. While this library is meant for dynamics, the SINDy techniques will work for non-temporal spatial problems by simply setting the time derivative equal to the forcing function, $\mathtt{x\_dot}[i]={\bm{f}}_{i}$.
The DeepONet is trained using the DeepXDE library (Lu et al., 2021b). The Neural Operator is trained using the neuralop library (Kossaifi et al., 2024).
The polynomial, sine, and cosine datasets were generated using SymPy by symbolically solving the differential equation for $u(x)$ for the form of $f(x)$. The polynomial data is generated using a root-form construction by $f(x)=\Pi^{p}_{i=1}(x-r_{i})$ for specified $p=[1,2,\ldots,8]$ with roots $r_{i}\sim U[-1,2]$. It is easier to control for ill-conditioning than with the monomial basis for high degree polynomials. The sine and cosine data are generated using the normal series in the text. The piecewise linear datasets were generated using the finite element method using FEniCS (Baratta et al., 2023).
Hyperparameters and training parameters for the models are given in Tables 1, 2, 3, 4, and 5.
Table 1: Hyperparameters for Linear model training (see Section 4.2)
Hyperparameter Values
Batch size Full batch training on 10k examples
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-1 to 1e-6
Epochs 2,000
Table 2: Hyperparameters for Deep Linear model training (see Section 4.2)
Hyperparameter Values
Layers 1
Hidden dimension 100
Batch size Full batch training on 10k examples
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-1 to 1e-6
Epochs 2,000
Table 3: Hyperparameters for MLP training (see Section 4.2)
Hyperparameter Values
Layers 1
Hidden dimension 1024
Activation function Leaky ReLU
Batch size 256
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-2 to 1e-6
Epochs 5,000
Table 4: Hyperparameters for Deep ONet training (see Section 4.2)
Hyperparameter Values
Layers 1
Branch x-tower dimensions [1, 256, 256]
Trunk F-tower dimensions [$N_{grid}$, 256, 256]
Activation function ReLU
Batch size 256
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Step from 1e-3 to 1e-4, 1e-5, 1e-6 every 5,000 steps
Training steps 20,000
Table 5: Hyperparameters for Fourier Neural Operator training (see Section 4.2)
Hyperparameter Values
Layers 1
N modes 64
Hidden dimension 32
Activation function GeLU
Batch size 256
Optimizer AdamW, $\lambda=0.01$
LR Scheduler Linear from 1e-2 to 1e-6
Epochs 5,000
@@ Appendix C Repeatability of test set generalization
Each run was repeated 5 times with different random seeds to verify the repeatability of the results. No seed results in a change in behavior for any for the models. The results are shown in Figures 7, 8, 9, 10, 11, and 12.
The finite difference model and linear model will always optimize to the same solution, where the linear model will have different noise in the null space. With the nonlinear models, it is possible that a lucky seed results in better generalization. However, we did not observe any. Repeated experiments with different seeds are shown in Figures 7, 8, 9, 10, 11, and 12. Each line on this plot corresponds to training on one function class identified in the legend, which corresponds to a row in the heat map from Figures 3 and 4. Only ten rows are shown for clarity. The error bars represent min/max over 5 runs, and the line is the average of the five. The repeated runs all show the same qualitative trends of generalization and overfitting in the same regimes for each of the six types of models. The repeated training runs are consistent with low variation in MSEs. The Neural Operator has the largest variability in MSEs between runs.
Figure 7: Generalization errors for the Finite Difference SINDy model with five different seeds for each training run. The finite difference optimization is invariant to the seed, so the min and max are within machine precision. The lower y limit is truncated to $10^{-12}$: training on Poly 1, Sine 1 and Cosine 1 achieves error $\approx 10^{-38}$.
Figure 8: Generalization errors for the linear model with five different seeds for each training run. With different random seeds, the min and max on the eval errors are within machine precision for the linear model.
Figure 9: Generalization errors for the deep linear model with five different seeds for each training run.
Figure 10: Generalization errors for the MLP with five different seeds for each training run.
Figure 11: Generalization errors for the DeepONet with five different seeds for each training run.
Figure 12: Generalization errors for the Neural Operator with five different seeds for each training run.
@@ Appendix D Grid size generalization
Figure 13: Convergence of the weight matrix $W$ as the grid size increases for two polynomial orders and piecewise linear forcing functions. The matrix structure is invariant to the discretization, however, each parameter scales proportional to $\Delta x$, as implied by Eq. 9.
Figure 14: Experiments on SINDy equation discovery as the grid size (x-axis) and training data polynomial function order (line color in legend) is changed. The number of parameters and equation libraries stays fixed as the number of points in the domain grows. 14(a): training data MSE (observable by the scientist). 14(b) : difference from true parameter (not observable). After convergence, the error tends to increase; we believe this is accumulation of floating point error as the points per example increases and not a fundamental limitation. I.e., with infinite precision we expect the error to go to 0.
While with time-dependent systems it is possible to measure $\Delta t$ between frames, we do not have a proposal for how to adapt models when $\Delta x$ in different examples may be variable and unknown. We consider the situation where we train and evaluate models on datasets produced at different grid spacings. (What if we bought camera A instead of camera B?) Note that $\Delta x$ is a property of the training data: decreasing $\Delta x$ in practice would require finer measurements by, for example, buying a better a camera with higher resolution.
In Fig. 13, we observe that the shape of the matrix does not change as the grid size increases: the model does not learn anything different. The polynomial datasets have the same number of coefficients for each grid size. However, the rank of “fem” dataset does increase as the piecewise linear functions are sampled the new grid.
For the finite difference method in Fig. 14, we do observe a difference in the error. As predicted by Theorem 2, the error on the learned parameter is proportional to both the grid size and polynomial order of the dataset. The figures also make it clear the order of the stencil does change the rate of convergence towards the true parameter. Increasing the order of accuracy of the finite difference approximation used is possible in practice, however, that increases the sensitivity to noise.

### v2

@@ A.2 Concrete Example of Error for 3 point stencil
For a given stencil, the above procedure can be executed algebraically to produced the graph in Fig. 2 (center, right). Plugging in the finite difference stencil into Equation 7, the inner value of the expression is
$\frac{w}{\Delta x^{2}}\Bigg(\left(\sum_{n=0}^{p+2}c_{n}(x-\Delta x)^{n}\right)-2\left(\sum_{n=0}^{p+2}c_{n}x^{n}\right)+\left(\sum_{n=0}^{p+2}c_{n}(x+\Delta x)^{n}\right)\Bigg)-k\sum_{n=2}^{p+2}n(n-1)c_{n}x^{n-2}$ (33)
For $p=0$ and $p=1$, the equation collapses to $w=k$. But when we consider $p=2$ with quadratic $f$ and quartic $u$, there is not full cancellation:
$\mathcal{L}=\mathbb{E}_{c}\Bigg[\int_{0}^{1}\Big(2w\left(c_{0}+3c_{1}x+6c_{2}x^{2}+c_{4}\Delta x^{2}\right)-2k\left(c_{0}+3c_{1}x+6c_{2}x^{2}\right)\Big)^{2}dx\Bigg]$ (34)
Following the procedure in the proof algebraically, we have
@@
which satisfies the general result and was used to plot the graph in Fig. 2 (center). A similar algebraic equation can be derived for the five point stencil to yield the graph in Fig. 2 (right).
A.3 Proof of Theorem 3.2
@@
A.4 Corollary to Theorem 3.2: Error Estimate
If the Green’s function and data basis is known, the analytical solution $\bm{A}$} can be computed to give estimate of error, shown in Fig.2. The procedure is as follows: Compute a matrix of basis functions evaluated at each point in the domain,
@@
${\bm{A}}_{ij}=\int_{\Omega}\psi_{i}(s)G(s,x_{j})\mathrm{d}x$ (46)
where $\psi_{i}$ is a piecewise linear or piecewise constant basis function centered on $x_{i}$. Then, the error starting at any initial matrix ${\bm{W}}^{0}$ can be directly computed as
$\|{\bm{W}}^{*}-{\bm{A}}\|_{F}\approx\|{\bm{W}}^{0}-{\bm{A}}{\bm{U}}{\bm{U}}^{T}\|_{F}$ (47)
where the computation for $\bm{A}$} is imperfect from the arbitrary choice of discretization basis $\psi(x)$.
@@ Appendix B Procedure Details
The ultimate experiments ran in float32 on a Macbook Pro M3 Max GPU using PyTorch. To initially rule out errors from accelerator hardware, drivers, or numerical precision, we replicated the linear models and deep models in many implementations: using Tinygrad, JAX, and PyTorch; on CPU, NVIDIA 4070 GPU, and Apple Silicon GPU hardware; and varied precision in float32 and float64. The behavior was consistent across all hardware and software variations.
[…]
All models were trained using AdamW with weight decay $\lambda=0.01$. The Linear and Deep Linear models used full-batch training on 10k examples for 2,000 epochs with a linear learning rate schedule from 1e-1 to 1e-6; the Deep Linear has one hidden layer of dimension 100. The MLP (1 layer, hidden dimension 1024, Leaky ReLU) and Fourier Neural Operator (64 modes, hidden dimension 32, GeLU) used batch size 256 for 5,000 epochs with a linear schedule from 1e-2 to 1e-6. The DeepONet used branch tower dimensions [1, 256, 256], trunk tower dimensions $N_{\text{grid}}$, 256, 256], ReLU activation, batch size 256, and a step schedule from 1e-3 decaying to 1e-6 every 5,000 steps over 20,000 total steps.
@@ Appendix C Repeatability of test set generalization
Repeated experiments with different seeds are shown in Figures 11(a), 11(b), 11(c), 11(d), 11(e), and 11(f). Each line on this plot corresponds to training on one function class identified in the legend, which corresponds to a row in the heat map from Figures 4 and 5. he error bars represent min/max over 5 runs, and the line is the average of the five. The repeated runs all show the same qualitative trends of generalization and overfitting in the same regimes for each of the six types of models.
(a) Finite Difference SINDy model.
(b) Linear model
(c) Deep linear model
(d) Multilayer Perceptron
(e) DeepONet
(f) Neural Operator
Figure 11: Generalization errors for all models with five different seeds for each training run. For the Finite Difference SINDy model, the lower y limit is truncated to $10^{-12}$; training on Poly 1, Sine 1 and Cosine 1 achieves error $\approx 10^{-38}$.
@@ Appendix D Grid size generalization
Figure 12: Convergence of the weight matrix $W$ as the grid size increases. The matrix structure is invariant to the discretization, however, each parameter scales proportional to $\Delta x$ (Eq. 9).
Figure 13: Experiments on SINDy equation discovery as the grid size (x-axis) and training data polynomial function order (line color in legend) is changed. 13(a): training data MSE (observable by the scientist). 13(b) : difference from true parameter (not observable). After convergence, the error tends to increase; we believe this is an artifact of accumulation of floating point error.
We consider the situation where we train and evaluate models on datasets produced at different grid spacings. Note that $\Delta x$ is a property of the training data: decreasing $\Delta x$ in practice would require finer measurements. In Fig. 12, we observe that the shape of the matrix does not change as the grid size increases. The polynomial datasets have the same number of coefficients for each grid size. However, the rank of “fem” dataset does increase as the piecewise linear functions are sampled the new grid. For the finite difference method in Fig. 13, we do observe a difference in the error. As predicted by Theorem 2, the error on the learned parameter is proportional to both the grid size and polynomial order of the dataset. The order of the stencil does change the rate of convergence towards the true parameter.